# Can black holes preserve $\mathcal{N}>4$ supersymmetry?

### v3

@@ 4.2 Commutation relations of conserved charges
[…]
@@ 4.4 Linearization of the model, its solution, and the index
$j_{-}+\frac{1}{2}$$j_{-}$$j_{-}-\frac{1}{2}$$j_{-}-1$$j_{-}-\frac{3}{2}$$j_{+}-\frac{1}{2}$$j_{+}-1$$j_{+}-\frac{3}{2}$$j_{+}$$j_{+}+\frac{1}{2}$  $j_{-}+\frac{1}{2}$$j_{-}$$j_{-}-\frac{1}{2}$$j_{-}-1$$j_{+}$$j_{+}-\frac{1}{2}$$j_{+}-1$$j_{+}+\frac{1}{2}$ (87)
where the white dots denote the states before the additional free fermions are incorporated. All states have the same $\text{U}(1)$ charge which we denote $u$. Another new aspect is that $\text{SU}(2)_{+}\otimes\text{SU}(2)_{-}$ representations have degeneracies. For example, for the long multiplet depicted in the left panel of the diagram above, there are $4$ representations with spins $(j_{+}-1/2,j_{-}-1/2)$, $2$ representations with spins $(j_{+}-1,j_{-}-1)$,$(j_{+}-1,j_{-})$, $(j_{+},j_{-}-1)$ and $(j_{+},j_{-})$, and the rest have only one. For the short multiplets all representations appear once except $(j_{+},j_{-}-1/2)$ and $(j_{+}-1/2,j_{-})$ which appear twice.
@@ 5 Cases with non-linearly realized symmetries
$I_{\text{on-shell}}=-\frac{2\pi^{2}\Phi_{r}}{\beta}-\frac{2\pi^{2}\Phi_{r}}{\beta}16\,\alpha_{\text{SU}(2)}^{2}+\frac{2\pi^{2}\Phi_{r}}{\beta}\sum_{i=1}^{n}8\,\alpha_{\text{Sp(2n)},i}^{2}$ (103)
@@
The second equation of the second line does not include the case $\alpha=-\beta$ and involves numerical factors $N_{\alpha\beta}$ fixed by Jacobi identities. The $12l+$ 1 supercharges lie in the fundamental representation of $\text{SO}(2l+1)$. We can choose the complex basis in which $Q^{\bar{i}}$ has the weight vector $e^{i}$, $Q^{i}$ has $-e^{i}$, and $Q^{0}$ has a vanishing weight vector; namely
@@
Here, $k,l$ run over $11,\dots 2l+$ 1 while $J^{a},J^{b}$ span all the generators in $\text{SO}(2l+1)$. Each generator $J^{a}$ is represented by $t^{a}_{ij}$ in the fundamental representation. The tensor $\Pi^{ab}_{kl}$ is defined as $\Pi_{ij}^{ab}=t_{im}^{a}t_{mj}^{b}+t_{im}^{b}t_{mj}^{a}+2\delta^{ab}\delta_{ij}$. The constants $c,K,\gamma$ can be determined through embedding in the $\text{SO}(n)$-extended Virasoro algebra Knizhnik:1986wc ; Bershadsky:1986ms . However, we leave them free here to be matched later from our spectrum. Among these anticommutators, we are particularly interested in the one involving zero weight supercharges:
@@
$Q^{\bar{i}}{\left|{E,\lambda}\right>}=$ $0\displaystyle 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}J^{\alpha}{\left|{E,\lambda}\right>}=$ 0 (117)
@@
Can these states preserve more supersymmetries for certain values of $m_{1},\dots,m_{l-1}\in\mathbb{Z}_{\geq 0}$? We can verify this by examining similar commutators $[J^{\alpha=-e^{i}},Q^{0}]=-\frac{1}{\sqrt{2}}Q^{i}$. In the nonzero directions of the weight vector, when $J^{\alpha=-e^{i}}$ acts on BPS states, it can generate states with weights $\lambda_{i}^{\prime}=\lambda_{i}-1$ in the $i$-th direction, making them deviate from the BPS energy. Therefore, from the algebra $\frac{1}{\sqrt{2}}Q^{i}{\left|{E_{\rm BPS},\lambda}\right>}=Q^{0}J^{\alpha=-e^{i}}{\left|{E_{\rm BPS},\lambda}\right>}\neq 0$ we can conclude that for these $i$, $Q^{i}$ cannot annihilate the BPS states. On the other hand, if $m_{l},m_{l-1},\dots,m_{l-k}$ are all zero, we can deduce from the same argument that these BPS states preserve $12(k+1)+$ 1 supersymmetries. The energy of these BPS states is still given by (119). As an example, when all $m$’s are zero and $k=l-1$, the states are fully BPS and in a singlet with energy $E_{\rm BPS}=\frac{(l-1)l(2l-1)}{48\Phi_{r}}$.
To summarize, most BPS short multiplets in this $\text{OSp}(2l+1|2)$ theory preserve three supercharges. However, setting certain charges to zero can make these multiplets even shorter and preserve additional supercharges ($12l+$ 1 in total at most), as discussed above. This is analogous to the distinction between standard BPS multiplets and shorter BPS multiplets (61) in the large $\mathcal{N}=4$ theory. This structure holds universally for all the following cases, but we will only describe the supersymmetry that general BPS states can preserve there.
@@
From the superalgebra isomorphism $D(2,1|1)\cong\text{OSp}(4|2)$, we expect the spectrum of these two theories should be identical. Actually, the BPS condition (67) for $D(2,1|\upalpha)$ becomes $j_{+}=j_{-}$ for $\upalpha=1$. Then we can match the spectrums of the two theories by identifying $\alpha_{-}=\alpha_{1}-\alpha_{2},\alpha_{+}=\alpha_{1}+\alpha_{2}$; $j_{+},j_{-}=\frac{m_{1}+1}{2},\frac{m_{2}}{2}$ in the continuous spectrum, and $j=\frac{m}{2}$ in the discrete spectrum. However, the restriction to integer-spin representations, $m_{1}+m_{2}\in 2\mathbb{Z}_{\geq 0}$, does not apply to $D(2,1|\upalpha)$, since its $R$-symmetry $\times\text{SU}(2)$ SU ( 2 ) is a double cover of $\text{SO}(4)$.
@@
  • The first has a $R$symmetry $\text{Spin}(n)$. This assumes the presence of fields in spinorial representations and, therefore, we need to sum over even shifts of the chemical potentials $n_{i}\in 2\mathbb{Z}$ in the partition function (19) and (20), instead of $n_{i}\in\mathbb{Z}$. The Laplace transform has multiplets with all representations, with no constraint on $m_{l}$ ($m_{l}+m_{l-1}$) being even for $n=2l+1$ odd ($n=2l$ even). Other than this, the expressions we found the density of states above still hold.
@@ 5.4 Exceptional superconformal groups
The long multiplet structure is analogous to the $\text{OSp}(2l+1|2)$ case: the overall factor of 2 comes from the action of a supercharge that is neutral under the $R$-symmetry; the action of three ‘lowering’ supercharge generates the multiplet of $82^{3}=$ 8 states, making the total dimension $82^{4}=$ 8 The explicit form of each character can be found in (LABEL:eqn:G2character). The density of long multiplets is given by
@@
In this subsection, we explore another example with ${\cal N}=8$ supersymmetry, the $\text{F}(4)$ Schwarzian theory. The eight fermions transform under the spinor representation of the $R$-symmetry $\text{Spin}(7)$, therefore have weights (±$(\pm\frac{1}{2},\pm\frac{1}{2},\pm\frac{1}{2})$. Given this information, along with the fact that $q=\frac{3}{4}$ for $\text{F}(4)$, we can write down the partition function as
@@
[…]
×$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((1+m_{1})^{2}+(2+m_{1}+m_{2})^{2}+(1+m_{2})^{2}\right)}$
×$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((1+m_{1})^{2}+(2+m_{1}+m_{2}+m_{3})^{2}+(1+m_{2}+m_{3})^{2}\right)}$
×$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((3+m_{1}+2m_{2}+m_{3})^{2}+(2+m_{1}+m_{2})^{2}+(1+m_{2}+m_{3})^{2}\right)}$
×$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((3+m_{1}+2m_{2}+m_{3})^{2}+(2+m_{1}+m_{2}+m_{3})^{2}+(1+m_{2})^{2}\right)}$
@@ Appendix A Technical details on the $\mathcal{N}=3$ and $\mathcal{N}=4$ theories
$-\partial_{\beta}\tilde{Z}=e^{S_{0}}2\cos 2\pi\alpha\sum_{n\in\mathbb{Z}}\frac{\alpha+n}{\sin 2\pi\alpha}\frac{8\pi\Phi_{r}}{\beta^{2}}\,e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-16(\alpha+n)^{2})}.$ (150)
@@
$\tilde{Z}$ $=$ $\sum_{j=\frac{1}{2},1,\ldots}\int_{E_{0}(j)}^{\infty}{\rm d}E(\chi_{j}(\alpha)+\chi_{j-1}(\alpha))\,\rho_{j}(E)e^{-\beta E}+\text{constant.}$. (154)
where the density of states is essentially the same as in $-\partial_{\beta}\tilde{Z}$ up to an overall factor of energy in the denominator, reproducing the formula quoted in the main text.
@@
In this case the derivation is more complex since the $R$-symmetry group of $D(2,1|\upalpha)$, $\times\text{SU}(2)$ SU ( 2 ), has rank 2. The partition function can be written as
@@
×$\quad\times\frac{e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left[1-\frac{4}{\gamma_{-}}\left(n+\alpha_{+}\right)^{2}-\frac{4}{\gamma_{+}}\left(m+\alpha_{-}\right)^{2}\right]}}{\left(1-4(\alpha_{+}+n-\alpha_{-}-m)^{2}\right)\left(1-4(\alpha_{+}+n+\alpha_{-}+m)^{2}\right)}$
@@
Here we introduce another sum over $\epsilon_{\pm}=\pm 1$ to emphasize $\times\text{SU}(2)$ SU ( 2 ) Weyl invariance, which acts by $\alpha_{i}\to-\alpha_{i}$. The individual saddle is given by
@@
×$\times\frac{e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left[1-\frac{4}{\gamma_{-}}(\alpha_{+}+n)^{2}-\frac{4}{\gamma_{+}}(\alpha_{-}+m)^{2}\right]}}{\left(\frac{1}{2}+\epsilon_{+}(\alpha_{+}+n)+\epsilon_{-}(\alpha_{-}+m)\right)}$
@@
×$\quad\times\cos\pi(n+m)\cos\pi(n-m)\frac{e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left(1-\frac{4}{\gamma_{-}}n^{2}-\frac{4}{\gamma_{+}}m^{2}\right)}}{\frac{1}{2}+n+m}$
@@
×$\times e^{\frac{2\pi^{2}\Phi_{r}}{\beta}-\frac{\beta(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int\frac{-{\rm d}t_{1}}{1+2t_{1}}e^{-\frac{2\pi^{2}\Phi_{r}}{\beta}4t_{1}^{2}+2\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
×$\times\int{\rm d}Ee^{-\beta E}\delta(E-\frac{(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})})\int\frac{-{\rm d}t_{1}}{1+2t_{1}}e^{2\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
×$\times\int{\rm d}Ee^{-\beta E}\delta(E-\frac{(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})})\frac{\pi}{2}e^{-\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}}\Theta(\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1})$
@@
×$\times\int{\rm d}Ee^{-\beta E}\delta\left(E-\frac{\upalpha(j_{+}+j_{-}+\frac{1}{2})^{2}}{2\Phi_{r}(1+\upalpha)^{2}}\right)\frac{\sqrt{\upalpha\pi/2}}{32(1+\upalpha)}\left|\sin\left(\frac{2\pi\upalpha(j_{+}+j_{-}+\frac{1}{2})}{1+\upalpha}\right)\right|$
@@
×$\times e^{-\frac{\beta(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int\frac{-{\rm d}t_{1}}{1+2t_{1}}(e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-4t_{1}^{2})}-1)e^{2\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
×$\times e^{-\frac{\beta(k_{+}-k_{-})^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int\frac{-{\rm d}t_{1}}{1+2t_{1}}(e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-4t_{1}^{2})}-1)e^{2\pi{\rm i}\frac{\upalpha k_{+}+k_{-}}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
[…]

### v4

@@ 4.2 Commutation relations of conserved charges
with $\updelta$, $\upbeta$ and $\upgamma$ arbitrary coefficients. Imposing Jacobi identities between the supercharges $\{Q_{A\dot{A}},Q_{B\dot{B}}\},Q_{C\dot{C}}]+\text{(cyclic)}=0$ leaves only one free parameter since it constrains $\updelta=\upbeta=\upgamma$. The final version of the algebra is
@@ 4.4 Linearization of the model, its solution, and the index
$_{-}+\frac{1}{2}jj_{-}$$j_{-}-\frac{1}{2}$$j_{-}-1$$j_{-}-\frac{3}{2}$$j_{+}-\frac{1}{2}$$j_{+}-1$$_{+}-\frac{3}{2}j+j_{+}$$j_{+}+\frac{1}{2}$  $_{-}+\frac{1}{2}jj_{-}$$j_{-}-\frac{1}{2}$$_{-}-1j+j_{+}$$j_{+}-\frac{1}{2}$$j_{+}-1$$j_{+}+\frac{1}{2}$ (87)
where the white dots denote the states before the additional free fermions are incorporated. All states have the same $\text{U}(1)$ charge which we denote $u$. Another new aspect is that $\text{SU}(2)_{+}\otimes\text{SU}(2)_{-}$ representations have degeneracies. For example, for the long multiplet depicted in the left panel of the diagram above, there are $4$ representations with spins $(j_{+}-1/2,j_{-}-1/2)$, $2$ representations with spins $(j_{+}-1,j_{-}-1)$$(j_{+}-1,j_{-})$, $(j_{+},j_{-}-1)$ and $(j_{+},j_{-})$, and the rest have only one. For the short multiplets all representations appear once except $(j_{+},j_{-}-1/2)$ and $(j_{+}-1/2,j_{-})$ which appear twice.
@@ 5 Cases with non-linearly realized symmetries
[…]
@@
The second equation of the second line does not include the case $\alpha=-\beta$ and involves numerical factors $N_{\alpha\beta}$ fixed by Jacobi identities. The $2l+1$ supercharges lie in the fundamental representation of $\text{SO}(2l+1)$. We can choose the complex basis in which $Q^{\bar{i}}$ has the weight vector $e^{i}$, $Q^{i}$ has $e^{i}$, and $Q^{0}$ has a vanishing weight vector; namely
@@
Here, $k,l$ run over $1,\dots 2l+1$ while $J^{a},J^{b}$ span all the generators in $\text{SO}(2l+1)$. Each generator $J^{a}$ is represented by $t^{a}_{ij}$ in the fundamental representation. The tensor $\Pi^{ab}_{kl}$ is defined as $\Pi_{ij}^{ab}=t_{im}^{a}t_{mj}^{b}+t_{im}^{b}t_{mj}^{a}+2\delta^{ab}\delta_{ij}$. The constants $c,K,\gamma$ can be determined through embedding in the $\text{SO}(n)$-extended Virasoro algebra Knizhnik:1986wc ; Bershadsky:1986ms . However, we leave them free here to be matched later from our spectrum. Among these anticommutators, we are particularly interested in the one involving zero weight supercharges:
@@
$Q^{\bar{i}}{\left|{E,\lambda}\right>}=$ $0,~~~~~~~~~~~J^{\alpha}{\left|{E,\lambda}\right>}=0$ (117)
@@
Can these states preserve more supersymmetries for certain values of $m_{1},\dots,m_{l-1}\in\mathbb{Z}_{\geq 0}$? We can verify this by examining similar commutators $J^{\alpha=-e^{i}},Q^{0}]=-\frac{1}{\sqrt{2}}Q^{i}$. In the nonzero directions of the weight vector, when $J^{\alpha=-e^{i}}$ acts on BPS states, it can generate states with weights $\lambda_{i}^{\prime}=\lambda_{i}-1$ in the $i$-th direction, making them deviate from the BPS energy. Therefore, from the algebra $\frac{1}{\sqrt{2}}Q^{i}{\left|{E_{\rm BPS},\lambda}\right>}=Q^{0}J^{\alpha=-e^{i}}{\left|{E_{\rm BPS},\lambda}\right>}\neq 0$ we can conclude that for these $i$, $Q^{i}$ cannot annihilate the BPS states. On the other hand, if $m_{l},m_{l-1},\dots,m_{l-k}$ are all zero, we can deduce from the same argument that these BPS states preserve $2(k+1)+1$ supersymmetries. The energy of these BPS states is still given by (119). As an example, when all $m$’s are zero and k=l1k=l-1, the states are fully BPS and in a singlet with energy $E_{\rm BPS}=\frac{(l-1)l(2l-1)}{48\Phi_{r}}$.
To summarize, most BPS short multiplets in this $\text{OSp}(2l+1|2)$ theory preserve three supercharges. However, setting certain charges to zero can make these multiplets even shorter and preserve additional supercharges ($2l+1$ in total at most), as discussed above. This is analogous to the distinction between standard BPS multiplets and shorter BPS multiplets (61) in the large $\mathcal{N}=4$ theory. This structure holds universally for all the following cases, but we will only describe the supersymmetry that general BPS states can preserve there.
@@
From the superalgebra isomorphism $D(2,1|1)\cong\text{OSp}(4|2)$, we expect the spectrum of these two theories should be identical. Actually, the BPS condition (67) for $D(2,1|\upalpha)$ becomes $j_{+}=j_{-}$ for $\upalpha=1$. Then we can match the spectrums of the two theories by identifying $\alpha_{-}=\alpha_{1}-\alpha_{2},\alpha_{+}=\alpha_{1}+\alpha_{2}$; $j_{+},j_{-}=\frac{m_{1}+1}{2},\frac{m_{2}}{2}$ in the continuous spectrum, and $j=\frac{m}{2}$ in the discrete spectrum. However, the restriction to integer-spin representations, $m_{1}+m_{2}\in 2\mathbb{Z}_{\geq 0}$, does not apply to $D(2,1|\upalpha)$, since its $R$-symmetry $\text{SU}(2)\times\text{SU}(2)$ is a double cover of $\text{SO}(4)$.
@@
  • The first has a $R$ -symmetry $\text{Spin}(n)$. This assumes the presence of fields in spinorial representations and, therefore, we need to sum over even shifts of the chemical potentials $n_{i}\in 2\mathbb{Z}$ in the partition function (19) and (20), instead of $n_{i}\in\mathbb{Z}$. The Laplace transform has multiplets with all representations, with no constraint on $m_{l}$ ($m_{l}+m_{l-1}$) being even for $n=2l+1$ odd ($n=2l$ even). Other than this, the expressions we found the density of states above still hold.
@@ 5.4 Exceptional superconformal groups
The long multiplet structure is analogous to the $\text{OSp}(2l+1|2)$ case: the overall factor of 2 comes from the action of a supercharge that is neutral under the $R$-symmetry; the action of three ‘lowering’ supercharge generates the multiplet of $2^{3}=8$ states, making the total dimension $2^{4}=8$ The explicit form of each character can be found in (LABEL:eqn:G2character). The density of long multiplets is given by
@@
In this subsection, we explore another example with ${\cal N}=8$ supersymmetry, the $\text{F}(4)$ Schwarzian theory. The eight fermions transform under the spinor representation of the $R$-symmetry $\text{Spin}(7)$, therefore have weights $(\pm\frac{1}{2},\pm\frac{1}{2},\pm\frac{1}{2})$. Given this information, along with the fact that $q=\frac{3}{4}$ for $\text{F}(4)$, we can write down the partition function as
@@
$(2+m_{1}+m_{2})(2+m_{1}+m_{2}+m_{3})(1+m_{2})(1+m_{2}+m_{3})\frac{\sinh\,2\pi\sqrt{2\Phi_{r}(E-E_{0})}}{E-E_{0}}$
$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((1+m_{1})^{2}+(2+m_{1}+m_{2})^{2}+(1+m_{2})^{2}\right)}$
$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((1+m_{1})^{2}+(2+m_{1}+m_{2}+m_{3})^{2}+(1+m_{2}+m_{3})^{2}\right)}$
$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((3+m_{1}+2m_{2}+m_{3})^{2}+(2+m_{1}+m_{2})^{2}+(1+m_{2}+m_{3})^{2}\right)}$
$\times\frac{1}{E-\frac{1}{18\Phi_{r}}\left((3+m_{1}+2m_{2}+m_{3})^{2}+(2+m_{1}+m_{2}+m_{3})^{2}+(1+m_{2})^{2}\right)}$
@@ Appendix A Technical details on the $\mathcal{N}=3$ and $\mathcal{N}=4$ theories
$\partial_{\beta}\tilde{Z}=e^{S_{0}}2\cos 2\pi\alpha\sum_{n\in\mathbb{Z}}\frac{\alpha+n}{\sin 2\pi\alpha}\frac{8\pi\Phi_{r}}{\beta^{2}}\,e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-16(\alpha+n)^{2})}.$ (150)
@@
$\tilde{Z}$ $=$ $\sum_{j=\frac{1}{2},1,\ldots}\int_{E_{0}(j)}^{\infty}{\rm d}E(\chi_{j}(\alpha)+\chi_{j-1}(\alpha))\,\rho_{j}(E)e^{-\beta E}+\text{constant.}$ (154)
where the density of states is essentially the same as in $\partial_{\beta}\tilde{Z}$ up to an overall factor of energy in the denominator, reproducing the formula quoted in the main text.
@@
In this case the derivation is more complex since the $R$-symmetry group of $D(2,1|\upalpha)$, $\text{SU}(2)\times\text{SU}(2)$, has rank 2. The partition function can be written as
@@
$\quad\times\frac{e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left[1-\frac{4}{\gamma_{-}}\left(n+\alpha_{+}\right)^{2}-\frac{4}{\gamma_{+}}\left(m+\alpha_{-}\right)^{2}\right]}}{\left(1-4(\alpha_{+}+n-\alpha_{-}-m)^{2}\right)\left(1-4(\alpha_{+}+n+\alpha_{-}+m)^{2}\right)}$
@@
Here we introduce another sum over $\epsilon_{\pm}=\pm 1$ to emphasize $\text{SU}(2)\times\text{SU}(2)$ Weyl invariance, which acts by $\alpha_{i}\to-\alpha_{i}$. The individual saddle is given by
@@
$\times\frac{e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left[1-\frac{4}{\gamma_{-}}(\alpha_{+}+n)^{2}-\frac{4}{\gamma_{+}}(\alpha_{-}+m)^{2}\right]}}{\left(\frac{1}{2}+\epsilon_{+}(\alpha_{+}+n)+\epsilon_{-}(\alpha_{-}+m)\right)}$
@@
$\quad\times\cos\pi(n+m)\cos\pi(n-m)\frac{e^{\frac{2\pi^{2}\Phi_{r}}{\beta}\left(1-\frac{4}{\gamma_{-}}n^{2}-\frac{4}{\gamma_{+}}m^{2}\right)}}{\frac{1}{2}+n+m}$
@@
$\times e^{\frac{2\pi^{2}\Phi_{r}}{\beta}-\frac{\beta(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int\frac{-{\rm d}t_{1}}{1+2t_{1}}e^{-\frac{2\pi^{2}\Phi_{r}}{\beta}4t_{1}^{2}+2\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
$\times\int{\rm d}Ee^{-\beta E}\delta(E-\frac{(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})})\int\frac{-{\rm d}t_{1}}{1+2t_{1}}e^{2\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
$\times\int{\rm d}Ee^{-\beta E}\delta(E-\frac{(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})})\frac{\pi}{2}e^{-\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}}\Theta(\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1})$
@@
$\times\int{\rm d}Ee^{-\beta E}\delta\left(E-\frac{\upalpha(j_{+}+j_{-}+\frac{1}{2})^{2}}{2\Phi_{r}(1+\upalpha)^{2}}\right)\frac{\sqrt{\upalpha\pi/2}}{32(1+\upalpha)}\left|\sin\left(\frac{2\pi\upalpha(j_{+}+j_{-}+\frac{1}{2})}{1+\upalpha}\right)\right|$
@@
$\times e^{-\frac{\beta(k_{+}-k_{-}+1)^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int\frac{-{\rm d}t_{1}}{1+2t_{1}}(e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-4t_{1}^{2})}-1)e^{2\pi{\rm i}\frac{\upalpha(k_{+}+\frac{1}{2})+(k_{-}-\frac{1}{2})}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
$\times e^{-\frac{\beta(k_{+}-k_{-})^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int\frac{-{\rm d}t_{1}}{1+2t_{1}}(e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-4t_{1}^{2})}-1)e^{2\pi{\rm i}\frac{\upalpha k_{+}+k_{-}}{\upalpha+1}t_{1}}\cos\pi t_{1}$
@@
$+e^{-\frac{\beta(k_{+}-k_{-})^{2}}{8\Phi_{r}(2+\upalpha+\upalpha^{-1})}}\int dt_{1}\frac{2\pi^{2}\Phi_{r}(1-2t_{1})}{\beta^{2}}e^{\frac{2\pi^{2}\Phi_{r}}{\beta}(1-4t_{1}^{2})+2\pi i\frac{\upalpha k_{+}+k_{-}}{\upalpha+1}t_{1}}\cos\pi t_{1}$